OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b, b2 − 4ac, 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

I really need help on this please!

OpenStudy (anonymous):

\[2x ^{2}+7x+3=0?\]

OpenStudy (anonymous):

so a=2 b=7 and c=3. does -b tell you if you can factor or not? what about the other components?

OpenStudy (anonymous):

basically the only rule that tells whether u can factor a quadratic into its roots is the discriminant \[\sqrt{b ^{2}-4ac}\]

OpenStudy (anonymous):

is the result is a perfect square, then you can complete the problem by factoring.

OpenStudy (anonymous):

if the discriminant is negative, then the roots are complex hope this helps.

OpenStudy (anonymous):

\[\sqrt{7^{2}-4(2)(3)}=\sqrt{49-24}=\sqrt{25}=5\]

OpenStudy (anonymous):

(2x +1)(x+ 3)=0 then solve for x \[x =-\frac{ 1 }{ 2 }\] or \[x = -3\]